OpenStudy (anonymous):

How do you find a gradient on a line with a 45 degree angle?

OpenStudy (anonymous):

you mean slope? it is 1

OpenStudy (anonymous):

yeah a slope, it's going from bottom left to top right and it has a 45 degree angle from the x axis

OpenStudy (anonymous):

i cant remember ;) but I think its 1

OpenStudy (anonymous):

but I have loads of questions on it and they can't all be 1 :/

OpenStudy (anonymous):

each question probably has a different theme or subject to it; what are the questions?

OpenStudy (anonymous):

the slope of a 45 degree angle; is 1

OpenStudy (anonymous):

the slope is simply the tangent; tan(45) = 1 right?

OpenStudy (anonymous):

how do I know whether to use tan sine or cos?

OpenStudy (amistre64):

sin tells you the y value cos tells you the x value and since slope = y/x ; slope = sin/cos = tan

OpenStudy (anonymous):

so that how you get the gradient?

OpenStudy (amistre64):

the gradient of an angle is sin(a)/cos(a) since sin/cos is how we define tan teh gradient of an angle = tan(a) tan(45) in this case = 1

OpenStudy (anonymous):

THANK YOU!!

OpenStudy (amistre64):

yep :)

OpenStudy (anonymous):

do you know how to calculate the size of the angle a straight line makes with a positive direction of the x axis. when you have the equation y=2x-3